hello i was wondering if anyone could help me with this word problem that i have been struggling with. it reads: find two whole numbers whose product is 15 such that one of the numbers is seven more than four times the other number. thank you for your time, it is greatly appreciated.
Answer by mananth(16946)

let the number be x& y
xy=15
x=15/y
..
x=4y+7
plug value of x
15/y=4y+7
15=y(4y+7)
15=4y^2+7y
4y^2+7y-15=0
4y^2+12y-5y-15=0
4y(y+3)-5(y+3)=0
(y+3)(4y-5)
so y =-3
...
plug value of y in the equation
x= -5
...
m.ananth@hotmail.ca